# Neural refractive index field: Unlocking the Potential of Background-oriented Schlieren Tomography in Volumetric Flow Visualization

### v1

@@ Abstract
keywords:
background-oriented schlieren, tomography, neural networks
1 Introduction
@@
Flow visualization techniques play a pivotal role in the field of fluid mechanics and combustion dynamics, offering detailed experimental data that is vital for understanding complex flow phenomena, enhancing control strategies, and verifying models in computational fluid dynamics (CFD) (Settles and Hargather, 2017). The advent of digital cameras and lasers has given rise to a multitude of visualization methods, each with its unique strengths: Particle Imaging Velocimetry (PIV) is renowned for its ability to create two-dimensional velocity maps (Lavoie et al., 2007), providing a snapshot of the flow velocity at a given plane; Planar Laser-Induced Fluorescence (PLIF) is adept at visualizing the distribution of minor chemical components in flames (Tanahashi et al., 2005), such as CH and OH, which are important intermediate species in combustion processes; and Filtered Rayleigh Scattering is used for temperature measurements (Carpenter and Pitz, 2020), offering insights into the thermal dynamics within the flow, to list a few.
@@
While these methods are significantly powerful, they often require expensive excitation lasers, complex optical systems, and specialized expertise for implementation. In contrast, Background-Oriented Schlieren (BOS) has gained widespread popularity since its introduction in 2000 due to its ease of implementation and versatility in measuring multiple physical parameters such as density, temperature and refractive index (Dalziel et al., 2000). It has been successfully applied in various challenging scenarios to visualize turbulent flows. For example, the solar disk has been used as a background for visualizing flow around a supersonic T-38 aircraft, captured by a ground-based imaging system equipped with a Calcium-K line filter (Hill et al., 2017). Similarly, BOS has been used to capture detailed schlieren images of a full-scale supersonic aircraft in flight, using the desert flora as a background (Heineck et al., 2021), resulting in the most detailed airborne schlieren imagery to date. Recently, the group of Grauer pioneered the integration of BOS with physics-informed neural networks (PINNs) and successfully demonstrated the method on axis-symmetric turbulent flows (Molnar et al., 2024, 2023a, 2023b). This approach further enhances the capability of BOS as it enforces the measured quantity to be consistent with the governing equations of the flows and allows for the inference of additional physical properties such as velocity and pressure.
@@
Despite the great success of these BOS methods, they are inherently two-dimensional and struggle to capture the full three-dimensional nature of turbulent flows (Grauer et al., 2023). Recently, Volumetric background-Oriented schlieren tomography (VBOST) has emerged as a promising technique (Raffel, 2015), which is widely recognized for its capability to retrieve the three-dimensional distribution of a spectrum of physical parameters (Liu et al., 2021). In recent years, significant progress has been made in VBOST, particularly in addressing measurement uncertainties arising from image displacement computations (Schmidt and Woike, 2021). These uncertainties are often attributed to the first-order Taylor approximation in optical flow algorithms and the discretization process in cross-correlation algorithms (Vinnichenko et al., 2023). Efforts to mitigate errors associated with optical flow have also been notable, with advancements in unified VBOST methods and hybrid approaches that combine voxel-based coarse reconstruction with light deflection considerations for fine reconstruction (Grauer and Steinberg, 2020).
@@
However, the existing VBOST modalities typically rely on voxel-based meshing scheme, which assumes constant physical properties within each voxel (Hu et al., 2024). This approach introduces several limitations, including significant discretization errors that result in severe reconstruction artefacts (Bo et al., 2023), exponential scaling of the weight matrix and memory requirements with increased voxel counts, and spatial resolution bottlenecked by a limited number of discrete voxels. To overcome these limitations, this work introduces a reconstruction approach termed Neural Refractive Index Field (NeRIF), which aims to train a light-weight neural network that receives the inputs of spatial coordinates of the points within the reconstruction domain and outputs both the corresponding refractive index and its gradient. The name of NeRIF draws inspiration from the cutting-edge computer vision algorithm called neural radiance field (NeRF) which has been straightforwardly adapted to tomographic techniques that reconstruct light radiance field  (Mildenhall et al., 2021; Zhang et al., 2023; Kelly and Thurow, 2023). Despite the conceptual parallel in utilizing neural networks, the physical models underpinning NeRF and NeRIF are inherently different. While the former focuses on modelling a radiance field, the latter is tailored to handle fields of refractive index and its gradient—highlighting a crucial divergence that requires a unique implementation strategy, marking a pivotal contribution of this research.
@@
To guarantee the superiority of NeRIF, three specially designed mechanisms are implemented. First, the reconstruction domain is implicitly represented with a compact neural network to avoid voxel-based discretization. Second, a tailored random sampling strategy is adopted to approximate the continuous integration and differentiation which are two essential steps for estimating the ray deflection, effectively mitigating discretization error as well as improving the spatial resolution. Third, both the output refractive index and the corresponding gradient are used to estimate the ray deflection for cross-validation, which prevents the neural networks from over-fitting the measurement noise. This step also involves the combination of the automatic and numerical differentiation mechanisms that helps expediting its convergence, thus improving the efficiency of the trained network (Chiu et al., 2022).
@@
2 Mathematical formulation of BOST
2.1 Voxel-based formulation
@@
2.2 Neural refractive index field
The algorithm here employs an implicit neural network-based approach (Mildenhall et al., 2021) to represent the 3D refractive index field spatially. Specifically, the target function of the refractive index and its corresponding gradient is formulated as $[\mathbf{n_{p}},\mathbf{\nabla n_{p}}]=f(\Theta,\mathbf{p})$, where $\mathbf{p}=(x,y,z)$ are input spatial coordinates, and $\Theta$ represents parameters of the neural network to be trained. Here, both the output refractive index and its gradient from this function are used separately to compute the corresponding sensor plane displacements based on the calibrated camera parameters. The calibration process is well-documented in (Grauer et al., 2023).
@@ 2.2 Neural refractive index field
We back-trace the rays from the camera pixel to the background board, as shown in Fig. 1(b). We can obtain a sequence of refractive indices and the gradients by inputting a series of sampling points along the ray path into the neural network. These values are then fed into Eq. 1 to calculate the displacements. The integration path theoretically extends from the camera to the background board. To minimize the computational cost, we define the integral range as $[t_{n},t_{f}]$, where $t_{n}$ and $t_{f}$ represent the distances of the farthest and the nearest ray points from the camera within the reconstruction domain. A random sampling strategy is then applied as follows: First, we set a random sampling number $N$, which follows a discrete uniform distribution $N\sim\text{Uniform}(M-10,M+10)$, where $M$ increases with the number of iterations, ranging from 60 to 200. Then, we divide the integration path into $N$ sampling intervals and randomly sample a point within each interval as
@@
where $t_{i}$ is the distance between the sampled point $\mathbf{p_{i}}$ (see Fig.1 (b)) and the camera pixel, and $U$ represents continuous uniform sampling. Finally, we apply numerical differentiation (ND) using the central difference method, with the step size $\Delta$ being selected as
@@
The constructed loss function is given in Eq. 7. It is crucial to acknowledge that when the coordinates of a sampled point are input into the mult-head neural network, it simultaneously and decoupledly generates both the refractive index and its gradients. However, a key aspect often overlooked is that these outputs are not inherently guaranteed to be consistent with each other. Specifically, if one were to take the derivative of the refractive index $(\nabla n)_{AD}$ output by the network via automatic differentiation (AD) and compare it with the directly predicted refractive index gradient $(\nabla n)_{\text{Net}}$, the two might not align perfectly. To address the mentioned inconsistency, we incorporate their difference into the loss function as the first term. In addition, both the refractive index and its gradients can be independently leveraged to predict displacements. Ideally, the predictions based on both quantities should converge on the same displacement values, matching those measured experimentally. To enforce this, we introduce two additional terms into the loss function. The first term measures the difference between the displacements $\hat{\varepsilon}_{i}$ predicted from the refractive index using discrete ND method and the experimentally measured displacements ${\varepsilon}_{i}$. Similarly, the second term quantifies the discrepancy between the displacements $\tilde{\varepsilon}_{i}$ predicted from the refractive index gradients and the same measured data. This approach integrates the mechanisms of ND and AD (Chiu et al., 2022), enhancing both the robustness and efficiency of the model, while ensuring accurate representation of the underlying physics.
$\text{Loss}=\frac{1}{N}\sum\left[\left\|\left(\nabla n\right)_{\text{AD}}-\left(\nabla n\right)_{\text{Net}}\right\|_{2}+\left\|\hat{\varepsilon}_{i}-\varepsilon_{i}\right\|_{2}+\left\|\tilde{\varepsilon}_{i}-\varepsilon_{i}\right\|_{2}\right].$ (7)
A fully connected neural network architecture with 7 layers of 256 neurons was adopted, incorporating a skip connection at the 4th layer, directly linking it to the input encoded with sinusoidal frequency encoding. The 7th layer connects to four individual single-layer neural networks, each with 64 outputs, to model $n$ and $\nabla n$ using a multi-head approach. The Adam optimizer was employed with a learning rate of $10^{-3}$ and decay rates of (0.9, 0.99).
3 Numerical and experimental validations
@@
The fourth-order Runge-Kutta method was utilized for refractive index fields to perform accurate ray tracing, resulting in the generation of nine simulated projection maps (each 200×200 pixels) at equidistant angular intervals spanning from 0 to 170 degrees. To mimic practical situations, varying levels of Gaussian noise were deliberately introduced into the simulated distorted maps. For comparative studies, the Tikhonov (Grauer et al., 2018) and Landweber (Liu et al., 2021) algorithms were employed without noise, each configured for a 40×40×40 voxel division. The simulation experiments were conducted on a high-performance server equipped with 64 cores of Intel® Xeon® Gold CPUs, 512 GB of RAM, and an NVIDIA RTX 4090 GPU, ensuring efficient and reliable computations.
@@
The experiments were conducted using a high-speed camera to capture the native flame with a temporal resolution of 1 kHz, which is fast enough to capture the evolution of the flame subjected to perturbation. The optical flow algorithm Deepflow  (Weinzaepfel et al., 2013) was then selected to infer the displacement vectors according to the comparison between algorithms(Grauer and Steinberg, 2020), etc.
4 Results
4.1 Numerical validation
@@
Figure 3(a) showcases the performance of NeRIF under various noise conditions in comparison to the Tikhonov and Landweber algorithms. For Phantom 1 featured with smooth density distribution, it proves challenging to discern differences in the reconstructions. However, in the case of Phantom 2, characterized by larger spatial density gradients and a more intricate, stochastic distribution, NeRIF outperforms the others by reconstructing a 3D density map with minimal visible errors. Conversely, the alternative algorithms exhibit notable reconstruction artefacts, highlighting the superiority of NeRIF in complex scenarios.
@@
Figure 3: (a) Reconstruction of 3D density field distribution with different noise levels and algorithms; (b) Reconstruction of density field distributions for slices at $z$= 40 mm under different noise levels, compared with two reference algorithms
@@
We undertake a further comparison of the algorithms against seven simulated spray flame scenarios. For each phantom, we analyze and evaluate the reconstructions of ten uniformly spaced slices, spanning a z-range from 5 to 75 mm, utilizing the Structural Similarity Index (SSIM) and correlation coefficient (CC) as assessment metrics. As depicted in Fig. 3(b), the voxel-based algorithm exhibits larger errors, with a CC hovering around 0.9, suggesting a moderate level of correlation with the phantoms. Furthermore, the SSIM values generally fall below 0.9, likely attributed to the inherent limitations of voxel discretization, which poses challenges in accurately reconstructing intricate details. In contrast, NeRIF achieves both SSIM and CC values that are approaching 1, underscoring its superior reconstruction quality. Notably, even as noise levels increase, the errors of NeRIF remain lower than those observed with voxel-based algorithms, demonstrating its robustness.
@@
To further quantify the reconstruction quality, the average reconstruction error is defined as $\left\|\mathbf{n}_{\text{gt}}-\mathbf{n}_{\text{rec}}\right\|_{2}/\left\|\mathbf{n}_{\text{gt}}-1\right\|_{2}$, termed the L2 error. Table 1 compares the algorithms across the entire reconstruction volume under varying noise levels. This comparison unequivocally demonstrates the superiority of NeRIF over voxel-based algorithms in terms of its robustness to noise and overall reconstruction quality. Notably, the CC for the reconstructed density fields using NeRIF consistently exceeds 0.99 across different noise levels, with L2 errors predominantly below 1%.
Table 1: Correlation Coefficients and $L2$ Errors for each Phantoms and Noise Levels
Ours Method(Noise Levels) Voxel-based Methods
Phantom Noise 0 Noise 0.05 Noise 0.1 Noise 0.15 Tikhonov Landweber
(corr / L2 error) (corr / L2 error) (corr / L2 error) (corr / L2 error) (corr / L2 error) (corr / L2 error)
Phamton1 0.9997 / 0.5848% 0.9998 / 0.4981% 0.9996 / 0.8169% 0.9992 / 0.8212% 0.9917 / 2.1395% 0.9931 / 1.8196%
Phamton2 0.9944 / 0.4229% 0.9942 / 0.4955% 0.9964 / 0.4190% 0.9958 / 0.3426% 0.6526 / 2.7613% 0.7599 / 2.3602%
Phamton3 0.9981 / 0.3936% 0.9980 / 0.3949% 0.9969 / 0.5032% 0.9971 / 0.6098% 0.8259 / 3.4058% 0.8809 / 2.8573%
Phamton4 0.9984 / 0.4854% 0.9981 / 0.4529% 0.9984 / 0.5903% 0.9984 / 0.4383% 0.8806 / 3.4767% 0.9155 / 2.9481%
Phamton5 0.9987 / 0.4557% 0.9983 / 0.7837% 0.9979 / 0.5419% 0.9981 / 0.5012% 0.8933 / 3.7339% 0.9256 / 3.1427%
Phamton6 0.9966 / 0.7831% 0.9980 / 0.9536% 0.9980 / 0.6393% 0.9973 / 0.8098% 0.9012 / 4.0234% 0.9315 / 3.3724%
Phamton7 0.9981 / 0.9062% 0.9985 / 0.5922% 0.9981 / 0.8367% 0.9983 / 1.0414% 0.9091 / 4.3036% 0.9370 / 3.6073%
@@
In contrast, NeRIF not only surpasses voxel-based methods in reconstruction accuracy but also achieves remarkable efficiency gains by significantly reducing computational resource consumption. Operating with approximately 10 GB of GPU memory, NeRIF’s compact reconstruction approach enables the spatial resolution of the reconstructed phantom to closely match that of the captured images, facilitating enhanced efficiency. As evidenced in Fig. 4(b), NeRIF leverages the parallel computing capabilities of GPUs to dramatically accelerate the reconstruction process, reducing the total time from over 10 hours to just over 20 minutes.

### v2

@@ Abstract
preprint: AIP/123-QED
Neural refractive index field: Unlocking the Potential of Background-oriented Schlieren Tomography in Volumetric Flow Visualization
I Introduction
@@
Flow visualization techniques play a pivotal role in the field of fluid mechanics and combustion dynamics, offering detailed experimental data that is vital for understanding complex flow phenomena, enhancing control strategies, and verifying models in computational fluid dynamics (CFD) (Settles & Hargather, 2017). The advent of digital cameras and lasers has given rise to a multitude of visualization methods, each with its unique strengths: Particle Imaging Velocimetry (PIV) is renowned for its ability to create two-dimensional velocity maps (Lavoie et al., 2007), providing a snapshot of the flow velocity at a given plane; Planar Laser-Induced Fluorescence (PLIF) is adept at visualizing the distribution of minor chemical components in flames (Tanahashi et al., 2005), such as CH and OH, which are important intermediate species in combustion processes; and Filtered Rayleigh Scattering is used for temperature measurements (Carpenter & Pitz, 2020), offering insights into the thermal dynamics within the flow, to list a few.
@@
While these methods are significantly powerful, they often require expensive excitation lasers, complex optical systems, and specialized expertise for implementation. In contrast, Background-Oriented Schlieren (BOS) has gained widespread popularity since its introduction in 2000 due to its ease of implementation and versatility in measuring multiple physical parameters such as density, temperature and refractive index (Dalziel et al., 2000; Raffel et al., 2000). It has been successfully applied in various challenging scenarios to visualize turbulent flows. For example, the solar disk has been used as a background for visualizing flow around a supersonic T-38 aircraft, captured by a ground-based imaging system equipped with a Calcium-K line filter (Hill & Haering Jr, 2017). Similarly, BOS has been used to capture detailed schlieren images of a full-scale supersonic aircraft in flight, using the desert flora as a background (Heineck et al., 2021), resulting in the most detailed airborne schlieren imagery to date. Recently, the group of Grauer and Cai pioneered the integration of BOS with physics-informed neural networks (PINNs) and successfully demonstrated the method on axis-symmetric turbulent flows (Molnar et al., 2024, 2023a, 2023b; Cai et al., 2021b, a). This approach further enhances the capability of BOS as it enforces the measured quantity to be consistent with the governing equations of the flows and allows for the inference of additional physical properties such as velocity and pressure.
@@
Despite the great success of these BOS methods, they are inherently two-dimensional and struggle to capture the full three-dimensional nature of turbulent flows (Grauer et al., 2023). Many teams have tried to solve the problem of three-dimensional flow field diagnosis, such as computed chemiluminescence tomography (CTC) (Cai et al., 2013; Liu et al., 2017) for measurement of chemical components, and tomographic particle image velocimetry (TPIV) for measurement of the three-dimensional velocity field (Agarwal et al., 2017). Recently, Volumetric background-oriented Schlieren tomography (VBOST) has emerged as a promising technique (Goldhahn & Seume, 2007; Raffel, 2015), which is widely recognized for its capability to retrieve the three-dimensional distribution of a spectrum of physical parameters (Liu et al., 2021, 2020). In recent years, significant progress has been made in VBOST, particularly in addressing measurement uncertainties arising from image displacement computations (Atcheson et al., 2009; Schmidt & Woike, 2021). These uncertainties are often attributed to the first-order Taylor approximation in optical flow algorithms and the discretization process in cross-correlation algorithms (Vinnichenko et al., 2023). Efforts to mitigate errors associated with optical flow have also been notable, with advancements in unified VBOST methods (Grauer & Steinberg, 2020) and hybrid approaches that combine voxel-based coarse reconstruction with light deflection considerations for fine reconstruction (Hu et al., 2024).
@@
However, the existing VBOST modalities typically rely on a voxel-based meshing scheme, which assumes constant physical properties within each voxel. This approach introduces several limitations, including significant discretization errors that result in severe reconstruction artifacts (Boas et al., 2012; wang2019high), exponential scaling of the weight matrix and memory requirements with increased voxel counts, and spatial resolution bottlenecked by a limited number of discrete voxels. To overcome these limitations, A lot of teams have made great efforts. Bo et al. employed a recurrent neural network based on Gated Recurrent Units (GRUs) to mitigate reconstruction artifacts generated during iterative processes to a certain extent. (Bo et al., 2023); Meanwhile, Li et al. utilized the visual hull (VH) method to decrease the number of voxels requiring reconstruction, effectively reducing artifacts while enhancing processing speed (li2024three). This work introduces a reconstruction approach termed Neural Refractive Index Field (NeRIF), which aims to train a light-weight neural network in the reconstruction process that receives the inputs of spatial coordinates of the points within the reconstruction domain and outputs both the corresponding refractive index and its gradient. The name of NeRIF draws inspiration from the cutting-edge computer vision algorithm called neural radiance field (NeRF) which has been straightforwardly adapted to tomographic techniques that reconstruct light radiance field  (Mildenhall et al., 2021; Zhang et al., 2023; Kelly & Thurow, 2023). Despite the conceptual parallel in utilizing neural networks, the physical models underpinning NeRF and NeRIF are inherently different. While the former focuses on modeling a radiance field, the latter is tailored to handle fields of refractive index and its gradient—highlighting a crucial divergence that requires a unique implementation strategy, marking a pivotal contribution of this research.
@@
To guarantee the superiority of NeRIF, three specially designed mechanisms are implemented. First, the reconstruction domain is implicitly represented with a compact neural network to avoid voxel-based discretization. Second, a tailored random sampling strategy is adopted to approximate the continuous integration and differentiation which are two essential steps for estimating the ray deflection, effectively mitigating discretization error as well as improving the spatial resolution. Third, both the output refractive index and the corresponding gradient are used to estimate the ray deflection for cross-validation, which prevents the neural networks from over-fitting the measurement noise. This step also involves the combination of the automatic and numerical differentiation mechanisms that helps expedite its convergence, thus improving the efficiency of reconstruction (Chiu et al., 2022).
@@
II Mathematical formulation of BOST
II.1 Voxel-based formulation
@@
II.2 Neural refractive index field
The specific implementation process of the algorithm is shown in Algorithm. 1. The algorithm here employs an implicit neural network-based approach (Mildenhall et al., 2021) to represent the 3D refractive index field spatially. Specifically, the target function of the refractive index and its corresponding gradient is formulated as $[\mathbf{n_{p}},\mathbf{\nabla n_{p}}]=f(\Theta,\mathbf{p})$, where $\mathbf{p}=(x,y,z)$ are input spatial coordinates, and $\Theta$ represents parameters of the neural network to be trained. Here, both the output refractive index and its gradient from this function are used separately to compute the corresponding sensor plane displacements based on the calibrated camera parameters. The calibration process is well-documented in (Grauer et al., 2023).
@@ II.2 Neural refractive index field
We back-trace the rays from the camera pixel to the background board, as shown in Fig. 1(b). We can obtain a sequence of refractive indices and gradients by inputting a series of sampling points along the ray path into the neural network. These values are then fed into Eq. 1 to calculate the displacements. The integration path theoretically extends from the camera to the background board. To minimize the computational cost, we define the integral range as $[t_{n},t_{f}]$, where $t_{n}$ and $t_{f}$ represent the distances of the farthest and the nearest ray points from the camera within the reconstruction domain. A random sampling strategy is then applied as follows: First, we set a random sampling number $N$, which follows a discrete uniform distribution $N\sim\text{Uniform}(M-10,M+10)$, where $M$ increases with the number of iterations, ranging from 60 to 200. Then, we divide the integration path into $N$ sampling intervals and randomly sample a point within each interval as
@@
where $t_{i}$ is the distance between the sampled point $\mathbf{p_{i}}$ (see Fig.1 (b)) and the camera pixel, and $U$ represents continuous uniform sampling. Compared with the voxel-based Monte Carlo sampling algorithm, the sampling interval on the backpropagation light corresponding to each pixel only contributes to the pixel itself, and random sampling ensures accuracy while greatly reducing the calculation amount.
Finally, we apply numerical differentiation (ND) using the central difference method, with the step size $\Delta$ being selected as
@@
The constructed loss function is given in Eq. 7. It is crucial to acknowledge that when the coordinates of a sampled point are input into the mult-head neural network, it simultaneously and decoupledly generates both the refractive index and its gradients. However, a key aspect often overlooked is that these outputs are not inherently guaranteed to be consistent with each other. Specifically, if one were to take the derivative of the refractive index $(\nabla n)_{AD}$ output by the network via automatic differentiation (AD) and compare it with the directly predicted refractive index gradient $(\nabla n)_{\text{Net}}$, the two might not align perfectly. We incorporate their difference into the loss function as the first term to address the mentioned inconsistency. In addition, both the refractive index and its gradients can be independently leveraged to predict displacements. Ideally, the predictions based on both quantities should converge on the same displacement values, matching those measured images experimentally. To enforce this, we introduce two additional terms into the loss function. The first term measures the difference between the integral displacements $\hat{\varepsilon}_{\text{ND}}$ predicted from the refractive index using the discrete ND method and the experimentally measured image displacements ${\varepsilon}$. Similarly, the second term quantifies the discrepancy between the integral displacements $\tilde{\varepsilon}$ calculated from the refractive index gradients output of NeRIF and the same measured data. This approach integrates the mechanisms of ND and AD (Chiu et al., 2022), enhancing both the robustness and efficiency of the model, while ensuring an accurate representation of the underlying physics.
$\text{Loss}=\frac{1}{N}\sum\left[\left\|\left(\nabla n\right)_{\text{AD}}-\left(\nabla n\right)_{\text{Net}}\right\|_{2}+\left\|\hat{\varepsilon}_{ND}-\varepsilon\right\|_{2}+\left\|\tilde{\varepsilon}-\varepsilon\right\|_{2}\right].$ (7)
A neural network architecture with 7 fully connected layers of 256 neurons was adopted, incorporating a skip connection at the 4th layer, directly linking it to the input encoded with sinusoidal frequency encoding. The 7th layer connects to four individual single-layer neural networks, each with 64 outputs, to model $n$ and $\nabla n$ using a multi-head approach. The Adam optimizer was employed with a learning rate of $10^{-3}$ and decay rates of (0.9, 0.99) (Diederik, 2014), and more detailed information is provided in Appendix. A.
1 Create a multi-head neural network and initialize
2 The displacements ${\varepsilon}$ of images from different views are obtained by optical flow algorithm
3 Trace rays in reverse from image coordinate points $(u,v)$ to obtain sampling points $p=(x_{w},y_{w},z_{w})$ in the world coordinate system as Eq. 5
4 while $\text{epoch}<n_{\text{outer}}$ and $\text{loss}>0.005\text{loss}_{\text{initial}}$ do
5       input p to neural network and output $n$ and $(\nabla n)_{\text{Net}}$
6       calculate the first term $\left\|\left(\nabla n\right)_{\text{AD}}-\left(\nabla n\right)_{\text{Net}}\right\|_{2}$ of the loss function through AD of $n$
7       calculate $\hat{\varepsilon}_{\text{ND}}$ and $\tilde{\varepsilon}$ through Eq. 1 and Eq. 6
8       backward the total loss as Eq. 7
9       epoch $\leftarrow$ epoch + 1
10
11The final density $\rho$ is derived from the G-D constant as Eq. 2
Algorithm 1 NeRIF
III Numerical and experimental validations
@@
Table 1: Numerical validation settings.
Simulated camera system
Pixel size 20 $\mu$m $\times$ 20 $\mu$m
Pixel resolution 200 $\times$ 200
Focal length 50 mm
Number of views 9
Angle $\theta$ $0^{\circ}-170^{\circ}$
Other settings
Volume of interest 50 mm $\times$ 50 mm $\times$ 50 mm
Noise type Gaussian
The fourth-order Runge-Kutta method was utilized for refractive index fields to perform accurate ray tracing, resulting in the generation of nine simulated projection maps (each 200$\times$200 pixels) at equidistant angular intervals spanning from 0 to 170 degrees, details of the implementation of numerical calculations are in Appendix. B. To mimic practical situations, varying levels of Gaussian noise were deliberately introduced into the simulated distorted maps. For comparative studies, the Tikhonov (Grauer et al., 2018) and Landweber (Liu et al., 2021) algorithms were employed without noise, each configured for a 40$\times$40$\times$40 voxel division. The simulation experiments were conducted on a high-performance server equipped with 64 cores of Intel® Xeon® Gold CPUs, 512 GB of RAM, and an NVIDIA RTX 4090 GPU, ensuring efficient and reliable computations. Detailed information on the simulation study is shown in Table. 1.
@@
The experiments were conducted using a high-speed camera to capture the native flame with a temporal resolution of 1 kHz, which is fast enough to capture the evolution of the flame subjected to perturbation. The Calibration of internal and external parameters for high-speed cameras is carried out by Matlab Camera Calibration Toolbox (MathWorks, 2023). The optical flow algorithm Deepflow  (Weinzaepfel et al., 2013) was then selected to infer the displacement vectors according to the comparison between algorithms(Grauer & Steinberg, 2020), etc. Detailed information on the experimental study is shown in Table. 2 and Table. 3.
Table 2: Specifications of the experimental system.
Camera
Exposure time 1 ms
Frame rate 1 kHz
Pixel resolution 1024 $\times$ 1024
Pixel size 20 $\mu$m $\times$ 20 $\mu$m
Camera Lens 50 mm and f/2.8
Objective lenses 24 mm and f/2.8
Fiber bundles
Resolution of fibers of input ends 355 $\times$ 420
Dimensionality of the input ends 5.2 mm $\times$ 5.2 mm
Dimensionality of the output end 16 mm $\times$ 16 mm
Table 3: Flow rates of various gases
Gas CH4 O2 N2
Flow Rate (SLPM) 1.00 7.93 2.00
IV Results
IV.1 Numerical validation
@@
Fig. 3 showcases the performance of NeRIF under various noise conditions in comparison to the Tikhonov and Landweber algorithms. For Phantom 1 featured with smooth density distribution, it proves challenging to discern differences in the reconstructions. However, in the case of Phantom 2, characterized by larger spatial density gradients and a more intricate, stochastic distribution, NeRIF outperforms the others by reconstructing a 3D density map with minimal visible errors. Conversely, the alternative algorithms exhibit notable reconstruction artifacts, highlighting the superiority of NeRIF in complex scenarios.In the slice graph Fig. 4 at z=40mm, we can intuitively see the advantages of NeRIF in reducing reconstruction artifacts.
@@
Figure 3: Reconstruction of 3D density field distribution with different noise levels and algorithms
Figure 4: 2D slices of reconstructed density field distribution with different noise levels and algorithms
@@
We undertake a further comparison of the algorithms against seven simulated spray flame scenarios. For each phantom, we analyze and evaluate the reconstructions of ten uniformly spaced slices, spanning a z-range from 5 to 75 mm, utilizing the Structural Similarity Index (SSIM) and correlation coefficient (CC) as assessment metrics, and detailed information is provided in Appendix. C. As depicted in Fig. 5, the voxel-based algorithm exhibits larger errors, with a CC hovering around 0.9, suggesting a moderate level of correlation with the phantoms. Furthermore, the SSIM values generally fall below 0.9, likely attributed to the inherent limitations of voxel discretization, which poses challenges in accurately reconstructing intricate details. In contrast, NeRIF achieves both SSIM and CC values that are approaching 1, underscoring its superior reconstruction quality. Notably, even as noise levels increase, the errors of NeRIF remain lower than those observed with voxel-based algorithms, demonstrating its robustness.
Figure 5: Structural similarity and correlation coefficients for different slices of each phantom
@@
To further quantify the reconstruction quality, the average reconstruction error is defined as $\left\|\mathbf{n}_{\text{gt}}-\mathbf{n}_{\text{rec}}\right\|_{2}/\left\|\mathbf{n}_{\text{gt}}-1\right\|_{2}$, termed the L2 error. Table. 4 and Table. 5 compares the algorithms across the entire reconstruction volume under varying noise levels. This comparison unequivocally demonstrates the superiority of NeRIF over voxel-based algorithms in terms of its robustness to noise and overall reconstruction quality. Notably, the CC for the reconstructed density fields using NeRIF consistently exceeds 0.99 across different noise levels, with L2 errors predominantly below 1%. We quantified the spatial resolution of the algorithm in Appendix. D.
@@
In contrast, NeRIF not only surpasses voxel-based methods in reconstruction accuracy but also achieves remarkable efficiency gains by significantly reducing computational resource consumption. Operating with approximately 10 GB of GPU memory, NeRIF’s compact reconstruction approach enables the spatial resolution of the reconstructed phantom to match that of the captured images closely, facilitating enhanced efficiency. As evidenced in Fig. 6(b), NeRIF leverages the parallel computing capabilities of GPUs to dramatically accelerate the reconstruction process, reducing the total time from over 10 hours to just over 20 minutes.
@@
Table 4: Correlation Coefficients for Various Phantoms and Noise Levels
Phantom Noise 0 Noise 0.1 Noise 0.15 Tikhonov Landweber
Phamton1 0.9997 0.9996 0.9992 0.9917 0.9931
Phamton2 0.9944 0.9964 0.9958 0.6526 0.7600
Phamton3 0.9981 0.9969 0.9971 0.8259 0.8809
Phamton4 0.9984 0.9984 0.9984 0.8806 0.9155
Phamton5 0.9987 0.9979 0.9982 0.8933 0.9256
Phamton6 0.9966 0.9980 0.9973 0.9012 0.9316
Phamton7 0.9981 0.9981 0.9983 0.9092 0.9370
Table 5: $L2$ Errors for Various Phantoms and Noise Levels
Phantom Noise 0 Noise 0.1 Noise 0.15 Tikhonov Landweber
Phamton1 0.585% 0.817% 0.821% 2.140% 1.820%
Phamton2 0.423% 0.419% 0.343% 2.761% 2.360%
Phamton3 0.394% 0.503% 0.610% 3.406% 2.857%
Phamton4 0.485% 0.590% 0.438% 3.477% 2.948%
Phamton5 0.456% 0.542% 0.501% 3.734% 3.143%
Phamton6 0.783% 0.639% 0.810% 4.023% 3.372%
Phamton7 0.906% 0.837% 1.041% 4.304% 3.607%
IV.2 Experimental validation
The numerical validation presented above has confirmed the robustness of the NeRIF technique. Now, we provide an experimental demonstration of this approach. Our study focused on a turbulent premixed methane/air flame with an equivalence ratio of $\phi=1.2$. Within the reactor’s central nozzle, depicted in Fig. 2, the fuel-air mixture was ignited amidst a nitrogen co-flow. Utilizing the NeRIF technique, we reconstructed the temporal evolution of the three-dimensional density field, as illustrated in Fig. 7(a). This reconstruction effectively circumvented the overlap of flame surface structures, enabling simultaneous visualization of the flame front leading edge and the interface between the hot combustion products and the nitrogen co-flow.
Given the absence of ground truth data for the direct comparison, we validated our reconstruction with an alternative approach. Out of the nine projections captured, eight were employed for the reconstruction process and one was saved for the validation. Fig. 7(b-c) showcase a comparison between the 9th measured projection and the re-projection simulated with the reconstructed refractive index field, in terms of metrics such as SSIM, PSNR, and CC. In order to explain the quality of the reconstructed image, we introduced the PSNR(Peak Signal to Noise Ratio), and detailed information is provided in Appendix. C.
For eight angles employed in the reconstruction process, the PSNR consistently surpasses 30 when comparing the measured projections with their corresponding re-projections across different phantoms. Furthermore, the SSIM exceeds 0.95, and the CC stands above 0.99, demonstrating a remarkable level of fidelity in the reconstructed three-dimensional density field. Taking the 0ms flame as an example, in Fig. 8 we visualized the internal structure of the reconstructed flame, and we can find that the unburned low-temperature region of the flame is well reconstructed, and has a clear flame interface. We also provide a video demonstration of the reconstructed 3D density field from 0 to 400 ms, available in Movie 1. As a result, the reconstruction can be considered as a highly reliable algorithm and utilized for further investigations on the intricate mechanism of turbulent flames. We also provide videos of Three-dimensional visualization of flame profiles with higher turbulence in Movie 2 and Movie 3.